A 25-ft by 35-ft rectangular swimming pool is surrounded by a walkway of uniform width. If the total area o the walkway is 396ft^2, how wide is the walkway?
Answer by mananth(16949)

A 25-ft by 35-ft rectangular swimming pool is surrounded by a walkway of uniform width. If the total area o the walkway is 396ft^2, how wide is the walkway?
let the width of walkaway be x
length of swimming pool + walkaway = 35+2x
width of swimming + walkaway = 25+2x
Total area = ( 35+2x)(25+2x)
total area - swimming pool area = walkaway area.
( 35+2x)(25+2x)- 25*35= 396
875+70x+50x+4x^2 -875= 396
70x+50x +4x^2=396
4x^2+120x-396=0
x^2+30x-99=0
x^2+33x-3x-99=0
x(x+33)-3(x+33)=0
(x-3)(x+33)=0
x=3 The width of the walaway = 3 feet